OpenStudy (anonymous):

If the domain is {-4, 0, 5}, find the range.

OpenStudy (jango_in_dtown):

what is the function?

OpenStudy (anonymous):

Given the function f(x) = 4 - 2x

OpenStudy (jango_in_dtown):

f(-4)=4-2(-4)=4+8=12 now calculate f(0) and f(5)

OpenStudy (jango_in_dtown):

f(0)=4-2(0)=4 and f(5)=4-2(5)=4-10=-6 so the range is {f(-4), f(0),f(5)}={12,4,-6}